Simplify combining like terms:
 3a – 2b – ab – (a – b + ab) + 3ab + b – a

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

When term have the same algebraic factors, they are like terms.
Then,
= 3a–2b–ab–a + b–ab + 3ab + b–a
= 3a–a–a–2b + b + b–ab–ab + 3ab
= a (3–1- 1) + b (-2 + 1 + 1) + ab (-1 -1 + 3)
= a (3- 2) + b (-2 + 2) + ab (-2 + 3)
= a (1) + b (0) + ab (1)
= a + ab

Thanks